# How to Leverage Predictive Uncertainty Estimates for Reducing Catastrophic Forgetting in Online Continual Learning

### v2

@@ 4. Methodology
In the following subsections, we describe the strategies for each of the two objectives, and the uncertainty scores considered in our evaluation. Then, we outline our uncertainty-aware memory management detailing the properties of the employed uncertainty estimate and its advantages compared to standard scores.
4.1. Memory Management
Similar to Chrysakis and Moens (2020), the memory population strategy is based on a class-balanced update, which is crucial to consider in case the stream of data is highly imbalanced. In fact, if classes are not equally represented in the memory, sampling from the memory may further deteriorate the predictive performance of the framework for under-represented classes. Differently from the populating strategy proposed in Chrysakis and Moens (2020), the criteria to decide which samples to keep in the memory is not random, but based on predictive uncertainty estimates in order to intentionally store samples with desired characteristics for each class. There are different ways to select the samples. We can decide to store a) the class-representative data points by selecting the least uncertain ones for each class (bottom-k); b) a diverse set of data points, spanning from the most representative to the most uncertain ones (step-size); and c) the easiest-to-forget samples by sampling the ones with high uncertainty (top-k).
From the second task on, we need a strategy for sampling from the memory a subset of data points used for replay (replay set). Since the memory represents a prototypical set of data points for each class, we assume that a random sampling is sufficient to extract informative data points from the memory. Following the standard practice, the number of samples in the replay set is equal to the batch size. It is important to note that the memory may already contain samples from the current task. In such cases, samples from the current task are excluded from the sampling process to ensure that the focus remains solely on those belonging to past tasks.
Figure 2. Illustration of the different aspects of uncertainty captured by model confidence scores and Bregman Information (BI). Close to the decision boundary we can notice how, due to the high aleatoric uncertainty, data points have low-confidence scores. Contrarily, due to high density of observed data, there is a low uncertainty about the data generating process resulting in a low BI.
4.2. Uncertainty in the Logit Space
Different measure of predictive uncertainty can capture distinct aspects of a model’s irreducible aleatoric uncertainty (inherent in the data) and its epistemic uncertainty (that stems from the finite training data and can be reduced by gathering more data). In online CL, the most commonly used measures are derived directly from the confidence scores of the model and mostly capture the irreducible aleatoric uncertainty (Wimmer et al., 2023). Here, we hypothesize it may be more beneficial for the model to replay instances which are representative in the sense that there is low uncertainty about the data generating process (we refer to this as low epistemic uncertainty, while acknowledging that there exist varying definitions). To not rely on specialized Bayesian models, we leverage a recently proposed bias-variance decomposition of the cross-entropy loss and compute a generalized variance term directly from the loss (Gruber and Buettner, 2023). Such bias-variance decomposition decomposes the expected prediction error (loss) into the model’s bias, variance, and an irreducible error (noise term). The latter is related to aleatoric uncertainty, whereas the variance term can directly be related to epistemic uncertainty (Gruber et al., 2023).
Gruber and Buettner (2023) have recently shown that a bias-variance decomposition of cross-entropy loss gives rise to the Bregman Information as the variance term and measures the variability of the prediction in the logit space. We illustrate the different aspects of uncertainty captured by confidence scores and BI respectively in Figure 2. For example, data points close to the decision boundary have a low confidence score due to the inherently high aleatoric uncertainty; in contrast, due to the high density of observed data, there is actually a low uncertainty about the data generating process (DGP), resulting in a low BI (low epistemic uncertainty). Outliers far away from the decision boundary can have a high confidence score, but will have a high BI due to the high uncertainty regarding the DGP. We hypothesize that it is samples with a low BI that are most useful for replay in online-FCL.
To populate the memory with representative samples, we therefore propose using an uncertainty estimator based on Bregman Information (BI) (Gruber and Buettner, 2023). The authors demonstrate that BI at the sample level can be estimated through deep ensembles or test-time augmentation (TTA). However, to reduce the computational overhead, we employ TTA for computing the estimations. Let us consider the problem of multi-class classification (as in our case), where the standard loss is represented by the cross-entropy. Considering a set $P$ of perturbations for a given data point $x$, we can compute the variance term of the classification loss $u(x)_{BI}$ as follows:
(1) $u(x)_{BI}=\frac{1}{P}\sum_{i=1}^{P}\text{LSE}(\hat{z}_{i})-\text{LSE}\left(\frac{1}{P}\sum_{i=1}^{p}\hat{z}_{i}\right),$
where $\hat{z}_{i}\in\mathbb{R}^{c}$ represent the logit predictions and $\text{LSE}(x_{1},\dots,x_{n})=\ln\sum_{i=1}^{n}e^{x_{i}}$ the LogSumExp (LSE) function respectively. Intuitively, a large value of $u(x)_{BI}$ means that the logits predicted across the perturbations vary significantly, suggesting a high uncertainty of the prediction and the DGP at this point of the input space. The use of this estimator is also motivated by the fact that, in comparison with other uncertainty scores such as entropy, smallest margin, or least confidence, there is no information loss in the estimation step. In fact, if we inspect these alternative metrics, we can notice that they either need a normalization step to move the logits in the probability space or rest on the largest activation value only. Furthermore, as reported in Gruber and Buettner (2023); Ovadia et al. (2019), and Tomani and Buettner (2021), common confidence scores are reliable only in case of well-calibrated models.
In contrast, the BI-based estimation of the epistemic uncertainty is meaningful also under distribution shift and able to identify robust and representative samples (Gruber and Buettner, 2023).
5. Experiments
5.1. Datasets and Settings
@@
Datasets. To understand the change in the performance when using different uncertainty strategies, we employ CIFAR-10 (Krizhevsky et al., 2009). To configure the online-CIL setup, we randomly assign with different random seeds a set of classes to 5 tasks. Thus, for each random seed, the class assignment per task is different. In this way, we can identify the strategy providing greater flexibility and effectiveness irrespective of the composition of the tasks. Once the best strategy is identified, we evaluate our findings on class-imbalanced scenarios to assess the performance under more realistic conditions. First, we employ two artificially controlled imbalanced datasets, CIFAR10-LT and CIFAR100-LT (Cao et al., 2019). As explained in Section 1, with this experiment we want to replicate a realistic scenario in which recent tasks contain less data than the older ones. For this, the most appropriate type of imbalance is the long-tailed (LT) one (Cui et al., 2019) which follows an exponential decay to choose the sample size of each class. Specifically, for both datasets, we decided to use an imbalance factor $\rho$ equal to $0.1$. This means that, if the largest class contains, e.g., $500$ data points, then the smallest one consists of $50$ instances. Finally, we decide to focus on biomedical image analysis using the microscopic peripheral blood cell images dataset (BloodCell) which consists of 8 classes annotated by expert clinical pathologists (Acevedo et al., 2020; Yang et al., 2023). To reflect the intended realistic conditions, we assign 2 classes for each tasks by increasing size. In this way, older tasks have more data points than the most recent ones.
Experimental Settings. The baseline for assessing the goodness of the proposed strategies will be Experience Replay (ER) (Chaudhry et al., 2019). Our decision to use ER is based on the fact that, despite its simplicity, it is surprisingly competitive when compared with more sophisticated and newer approaches as shown in recently conducted empirical surveys (Soutif-Cormerais et al., 2023).
@@
In all the experiments, we employ a slim version of Resnet18 (He et al., 2016) – as done in previous works (Hurtado et al., 2023; Lopez-Paz and Ranzato, 2017; Soutif-Cormerais et al., 2023; Kumari et al., 2022) –, and use the SGD optimizer with a learning rate of $0.1$. Following standard practice, we set the batch size equal to $10$. We set the memory size to different values for each dataset to evaluate a variety of memory configurations considering both large or small buffers. For evaluation, we employ last accuracy (A) and last forgetting (F) defined in Chaudhry et al. (2018a). The compared uncertainty scores are least confidence (LC), smallest margin (SM), ratio of confidence (RC), entropy (EN), and the rainbow memory agreement score (RM) (Shannon, 1948; Campbell et al., 2000; Lewis and Gale, 1994; Culotta and McCallum, 2005; Bang et al., 2021). All the experiments are run on three different random seeds.
Table 1. Comparison of last accuracy (A) and last forgetting (F) on CIFAR10.
@@
5.2. Empirical Results
Table 1 shows the results for all the uncertainty metrics and sorting strategies employed for populating the memory. From the values reported in the tables, we can observe that selecting the most representative samples for each class (’Bottom’ column) consistently improves the results in terms of both accuracy (A) and forgetting (F) across all the considered uncertainty scores and memory sizes. Thus, our results indicate that the easiest-to-remember samples are beneficial in improving the predictive performance of the learning model and in reducing considerably CF.

### v3

@@ 4 Methodology
In the following subsections, we describe the available strategies for each of the two objectives, and the uncertainty scores considered in our evaluation. Finally, we propose an alternative uncertainty estimate that overcomes the limitations of the most popular uncertainty metrics.
4.1 Memory Management
Memory Population.
Apart from the criteria-based population strategies presented in Section 2.2.1, random strategies have demonstrated to work surprisingly well in online-CL. Reservoir (Vitter, 1985) is a random sampling technique without replacement to select $n$ samples from a pool of $N$ samples where $N$ is unknown and $N>n$. This strategy is used in Chaudhry et al. (2019) as a way to populate a memory of size $M$ from a stream of data points with unknown length. In this way, each data point has a probability of being included in the memory equal to $\frac{M}{N}$. Class-balanced Reservoir, instead, is based on a class-balanced random sampling strategy (Chrysakis & Moens, 2020). This is important to consider the case where the stream of data is highly imbalanced. Indeed, class-imbalance may further deteriorate the predictive performance of the framework.
Different from the population strategy proposed in Chrysakis & Moens (2020), where each class-memory set is populated and updated randomly, we introduce a class-balanced memory management based on predictive uncertainty estimates. This is similar to Bang et al. (2021), where the intent is to keep in the memory samples with desired characteristics for each class.
Memory Sampling.
Starting from the second task, we need a strategy to sample data points from the memory. The easiest way is random sampling, which selects at random a subset from the available data points in the memory buffer excluding the ones from the current task. The size of the replay set is equal to the batch size. Another option would be to extract past samples with specific characteristics. For this purpose, we can exploit uncertainty estimates. However, this solution represents an effective strategy only in the case the memory is populated at random. Furthermore, given the size of the memory buffer and the frequency of memory updates, computing uncertainty estimates for the whole buffer may be computationally expensive and not feasible during training.
In our case, since the memory is populated with samples with specific characteristics, the random sampling is sufficient.
4.2 Uncertainty Metrics
4.2.1 Predictive Uncertainty Estimation
Following previous works (Bang et al., 2021; Wang et al., 2022), to compute predictive uncertainty estimations for a given input image $x$, we employ the ensembling technique Test-Time Augmentation (TTA) (Wang et al., 2019). The set of transformations applied are the ones usually employed for image classification. A detailed list of the transformations used in our evaluation can be found in the supplementary. After selecting the set of transformations, we apply them on the selected input. In this way, we create a set of $P$ perturbed inputs $\tilde{x}$ which are used at test-time to compute their corresponding logits $\hat{z}$. Finally, the generated logits are used by the uncertainty estimates in different ways to compute the uncertainty score $u(x)$.
4.2.2 Predictive Uncertainty Scores
In our investigation, we consider the following popular uncertainty metrics:
    Least Confidence (LC) (Culotta & McCallum, 2005) determines the level of predictive uncertainty by examining the samples with the smallest predicted class probability. A model is less certain about the sample if it associates the most probable class $y_{(1)}$ with a low probability.
    $u(x)_{LC}=1-\frac{1}{P}\sum_{i=1}^{P}p(y_{(1)}=\hat{c}|\tilde{x}_{i})$ (2)
    Margin Sampling (MS) (Campbell et al., 2000) calculates predictive uncertainty by computing the difference between the most probable predicted class $y_{(1)}$ and the second largest one $y_{(2)}$. If the difference is low, the model is uncertain.
    $u(x)_{MS}=1-\frac{1}{P}\sum_{i=1}^{P}\left(p(y_{(1)}=\hat{c}|\tilde{x}_{i})-p(y_{(2)}=\hat{c}|\tilde{x}_{i})\right)$ (3)
    Ratio of Confidence (RC) (Campbell et al., 2000) is similar to MS. In this case, the ratio between the probabilities of the two most probable classes is computed. The closer the ratio is to 1, the more uncertain the model is about the considered sample.
    $u(x)_{RC}=\frac{1}{P}\sum_{i=1}^{P}\left(\frac{p(y_{(2)}=\hat{c}|\tilde{x}_{i}))}{p(y_{(1)}=\hat{c}|\tilde{x}_{i}))}\right)$ (4)
    Entropy (EN) (Shannon, 1948), differently from the above scores, estimates the uncertainty considering the whole probability distribution. If the predicted probabilities are similar to each other (i.e., tend to an uniform distribution), the model is not certain about the sample. The formulation reads as follows:
    $u(x)_{EN}=-\frac{1}{P}\sum_{i=1}^{P}\left(\sum_{j}p(y_{j}=\hat{c}|\tilde{x}_{i})\log(p(y_{j}=\hat{c}|\tilde{x}_{i}))\right)$ (5)
In addition to these metrics, we will also consider the agreement score $u(x)_{RM}$ reported in Eq. (1).
4.2.3 Uncertainty Quantification via Bregman Decomposition
Inspecting the metrics described above, we can make two main observations: 1) all the metrics need a normalization step to squash the logits in the $[0,1]$ range of the probability space; 2) the majority of them compute the uncertainty relying on the largest predicted probabilities only. In both cases, with different magnitudes, there is a loss of information compared to using the logit space. Importantly, such confidence scores are a reliable measure of predictive uncertainty only when the model is well calibrated, which is not the case in many real-world scenarios (Gruber & Buettner, 2023; Ovadia et al., 2019; Tomani & Buettner, 2021).
Furthermore, the confidence scores detailed in Section 4.2.2 mostly capture the aleatoric uncertainty which is inherent in the data and irreducible (Wimmer et al., 2023). For this reason, we believe it would be favorable to focus on samples with low epistemic uncertainty which is model-centric and can be reduced by gathering more data. To provide an intuitive explanation behind this decision, let us consider Figure 2 where we have a graphical representation of the behaviour of the two different sources of uncertainty. With commonly used confidence scores, regions with low confidence correspond to areas close to the decision boundary irrespective of the presence or not of outliers. Contrarily, when exploiting measures mostly based on epistemic uncertainty, low confidence regions correspond to areas in which the data density is low, i.e., where the uncertainty could be reduced by gathering more data. Consequently, if we are interested in the high confidence samples, this would result in selecting different samples depending on the employed score. For confidence scores, given their distance from the decision boundary, we may obtain outliers not being representative of the class of interest.
In light of these considerations, following recent advances for computing uncertainty estimates in the logit space for classification tasks, we suggest to employ an uncertainty estimator based on the Bregman Information (BI) (Gruber & Buettner, 2023). The authors demonstrate that BI can be used to quantify the variance of the model at the sample level through deep ensembles or TTA. This approach is particularly appealing from a theoretical standpoint, since it stems from a general bias-variance decomposition of proper scoring rules which gives rise to the Bregman Information as the variance term of the loss.
For the commonly used cross-entropy loss, this variance term – that can be directly associated to the epistemic uncertainty – can be estimated via $P$ TTA perturbations as follows:
$u(x)_{BI}=\frac{1}{P}\sum_{i=1}^{P}\text{LSE}(\hat{z}_{i})-\text{LSE}\left(\frac{1}{P}\sum_{i=1}^{p}\hat{z}_{i}\right),$ (6)
where $\hat{z}_{i}\in\mathbb{R}^{c}$ and $\text{LSE}(x_{1},\dots,x_{n})=\ln\sum_{i=1}^{n}e^{x_{i}}$ represent the logits generated by the model and the LogSumExp (LSE) function respectively. Intuitively, a large value of $u(x)_{BI}$ means that the logits predicted across the perturbations are very different and, thus, the model is not confident about the considered sample.
Figure 2: Illustration of the behaviour of the Bregman Information for uncertainty estimation in classification tasks. Given the focus on epistemic uncertainty, BI is low in the presence of high data density areas – i.e., where the uncertainty could be reduced by gathering more data.
4.2.4 Uncertainty Scores Ranking
Depending on the characteristics we want to extract from the images, we can use the uncertainty estimates in different ways. In particular, in our experiments we consider three different strategies to select the set of images for populating the memory depending on the uncertainty scores. By sampling images with the highest uncertainty (top-k), we aim at extracting the easiest-to-forget samples which are closer to the decision boundaries and/or represent outliers. With the step-sized strategy, we aim at sampling a diverse set of images, spanning from the most representative to the most uncertain ones. Finally, with the bottom-k approach we aim at finding the most representative set of images for each of the classes seen in the past tasks. Note that, as explained above, different uncertainty scores select different images based on their greater focus on epistemic or aleatoric uncertainty.
5 Experiments
5.1 Datasets and Settings
5.1.1 Datasets
@@
To understand the change in the performance when using different uncertainty strategies, we employ two datasets commonly used in online-CL for image classification, CIFAR-10 and CIFAR-100 (Krizhevsky et al., 2009). To configure the online-CIL setup, we randomly assign with different random seeds a set of classes to 5 tasks. Thus, each task $T_{t}$ has 2 (CIFAR-10) or 20 (CIFAR-100) classes designated. In the second part of our experiments, once the best strategy is identified, we evaluate our findings on class-imbalanced scenarios to assess the performance under more realistic conditions. First, we employ two artificially controlled imbalanced datasets, CIFAR10-LT and CIFAR100-LT (Cao et al., 2019). As explained in Section 1, with this experiment we want to replicate a realistic scenario in which recent tasks contain less data than the older ones. For this, the most appropriate type of imbalance is the long-tailed (LT) one (Cui et al., 2019) which follows an exponential decay to choose the sample size of each class. Specifically, for both datasets, we decided to use an imbalance factor $\rho$ equal to $0.1$. This means that, if the largest class contains, e.g., $500$ data points, then the smallest one consists of $50$ instances. Finally, we decide to focus on biomedical image analysis using the microscopic peripheral blood cell images dataset (BloodCell) which consists of 8 classes annotated by expert clinical pathologists (Acevedo et al., 2020; Yang et al., 2023). To reflect the intended realistic conditions, we assign 2 classes for each tasks by increasing size. In this way, older tasks have more data points than the most recent ones. This simulates a realistic scenario where most common subtypes are prevalent when collecting data for a ML-model and, over time, less common disease subtypes appear in the clinic and the model needs to be updated with the new incoming classes. We provide more details about the motivations behind the choice of this dataset in Appendix A.4.
5.1.2 Experimental Settings
The baseline for assessing the goodness of the proposed strategies will be Experience Replay (ER) (Chaudhry et al., 2019). ER consists of a reservoir approach for the memory management part, and a random sampling for replay. We decide to use ER because, despite its simplicity, it is surprisingly competitive compared to more sophisticated and newer approaches. This is corroborated by recent empirical surveys finding that newly introduced methods perform very similarly to the common Experience Replay (ER) method (Soutif-Cormerais et al., 2023). In addition to the standard strategy, we will also consider the class-balanced version (CBR) proposed in Chrysakis & Moens (2020) and previously described in Section 4.1, and a gradient-based method (A-GEM Chaudhry et al. (2018b)) for comparison. Finally, we assess the results in comparison with Monte Carlo Dropout (MC) (Gal & Ghahramani, 2016). As reported in Ovadia et al. (2019), MC can be used to estimate predictive uncertainty and considered as a competitive baseline under distribution shifts. Since online-CL can be seen as a special case of data under distribution shifts, we include MC as an additional baseline to estimate predictive uncertainty from a Bayesian perspective.
@@ 5.1.2 Experimental Settings
In all the experiments, we employ a slim version of Resnet18 (He et al., 2016) – as done in previous works (Hurtado et al., 2023; Lopez-Paz & Ranzato, 2017; Soutif-Cormerais et al., 2023; Kumari et al., 2022) –, and use the SGD optimizer with a learning rate of $0.1$. Following standard practice, we set the batch size equal to $10$. We set the memory size to different values for each dataset to evaluate a variety of memory configurations considering both large or small buffers.
For evaluation, following previous works, we employ the Last Accuracy (A) and Last Forgetting (F) – defined in Chaudhry et al. (2018a). Last refers to the calculation of the metrics at the end of the training on all tasks. Let denote with $a^{t,i}$ and $T$, the accuracy of task $i$ after learning task $t$ and the total number of tasks respectively.
The last accuracy $A$ is then defined as follows:
$A=\frac{1}{T}\sum_{i=1}^{T}a^{T,i}.$ (7)
Last Forgetting (F) is defined as the difference between the peak knowledge (i.e., the maximum accuracy) captured about a particular task during the learning process and its accuracy at the end of the learning process. This provides an idea of the information retained about a certain task after learning new tasks. The last forgetting (F) is computed as:
$F=\frac{1}{T-1}\sum_{t=1}^{T-1}\max_{l\in\{1,\dots,T-1\}}a^{l,j}-a^{T,j},\qquad\forall j<T.$ (8)
All the experiments are run on three different random seeds. The code to conduct the experiments is available at https://github.com/MLO-lab/uncertainty_estimates_for_CF.
Table 1: Comparison of last accuracy (A) and last forgetting (F) on CIFAR10.
@@
Table 2: Comparison of last accuracy (A) and last forgetting (F) on CIFAR100.
Score M=1000 M=2000
ER 14.23 $\pm$ 1.00 16.84 $\pm$ 0.96
29.32 $\pm$ 0.15 22.67 $\pm$ 1.35
CBR 13.38 $\pm$ 0.58 16.59 $\pm$ 0.41
31.66 $\pm$ 0.30 23.05 $\pm$ 1.29
A-GEM 14.06 $\pm$ 0.46 16.52 $\pm$ 1.05
33.23 $\pm$ 1.81 28.80 $\pm$ 0.79
Top Step Bottom Top Step Bottom
LC A()($\uparrow$)( ↑ ) 13.33 $\pm$ 0.16 12.52 $\pm$ 0.65 14.07 $\pm$ 0.36 15.63 $\pm$ 1.28 15.06 $\pm$ 0.69 15.44 $\pm$ 1.31
F()($\downarrow$)( ↓ ) 25.18 $\pm$ 1.51 27.50 $\pm$ 1.02 22.67 $\pm$ 3.02 19.71 $\pm$ 2.82 20.56 $\pm$ 1.93 17.50 $\pm$ 3.48
MS A()($\uparrow$)( ↑ ) 13.69 $\pm$ 1.25 11.69 $\pm$ 1.07 13.34 $\pm$ 0.22 15.76 $\pm$ 0.93 14.73 $\pm$ 0.92 16.32 $\pm$ 1.11
F()($\downarrow$)( ↓ ) 24.45 $\pm$ 1.82 27.05 $\pm$ 1.20 25.64 $\pm$ 3.62 20.23 $\pm$ 2.76 18.14 $\pm$ 1.26 17.50 $\pm$ 3.81
RC A()($\uparrow$)( ↑ ) 13.65 $\pm$ 0.97 12.65 $\pm$ 0.07 13.94 $\pm$ 0.78 16.69 $\pm$ 1.77 14.30 $\pm$ 0.80 16.03 $\pm$ 0.88
F()($\downarrow$)( ↓ ) 25.71 $\pm$ 2.99 27.55 $\pm$ 0.48 24.99 $\pm$ 3.48 17.15 $\pm$ 3.25 20.72 $\pm$ 1.50 17.68 $\pm$ 2.86
EN A()($\uparrow$)( ↑ ) 13.00 $\pm$ 0.64 12.79 $\pm$ 0.30 13.42 $\pm$ 0.87 14.18 $\pm$ 1.77 14.49 $\pm$ 0.73 14.42 $\pm$ 0.78
F()($\downarrow$)( ↓ ) 25.30 $\pm$ 2.25 27.82 $\pm$ 1.80 23.20 $\pm$ 4.00 20.93 $\pm$ 3.83 19.59 $\pm$ 2.58 17.43 $\pm$ 2.61
RM A()($\uparrow$)( ↑ ) 12.99 $\pm$ 1.35 12.24 $\pm$ 0.22 13.65 $\pm$ 0.75 15.62 $\pm$ 0.99 14.33 $\pm$ 0.64 16.39 $\pm$ 0.52
F()($\downarrow$)( ↓ ) 26.01 $\pm$ 3.18 28.08 $\pm$ 0.98 24.27 $\pm$ 3.52 19.49 $\pm$ 4.25 19.25 $\pm$ 1.55 16.68 $\pm$ 3.67
MC A()($\uparrow$)( ↑ ) 9.48 $\pm$ 1.92 10.49 $\pm$ 1.05 10.77 $\pm$ 2.71 10.99 $\pm$ 1.16 12.36 $\pm$ 1.64 10.68 $\pm$ 2.75
F()($\downarrow$)( ↓ ) 16.77 $\pm$ 2.42 15.09 $\pm$ 1.41 7.50 $\pm$ 2.53 15.98 $\pm$ 1.45 11.91 $\pm$ 0.90 9.08 $\pm$ 4.36
BI A()($\uparrow$)( ↑ ) 13.38 $\pm$ 0.59 12.27 $\pm$ 0.69 13.81 $\pm$ 1.00 14.52 $\pm$ 0.54 14.33 $\pm$ 1.41 17.03 $\pm$ 0.30
F()($\downarrow$)( ↓ ) 25.43 $\pm$ 3.23 25.50 $\pm$ 1.15 19.41 $\pm$ 1.74 20.87 $\pm$ 3.51 17.68 $\pm$ 1.63 13.30 $\pm$ 3.65
5.2 Empirical Results
Tables 1 and 2 show the results for all the uncertainty metrics and sorting strategies employed for populating the memory. From the values reported in the tables, we can observe that selecting the most representative samples for each class (’Bottom’ column, in green) consistently improves the results in terms of both accuracy (A) and forgetting (F) across all the considered uncertainty scores and memory sizes. In some cases, ER is able to outperform our proposed approach in terms of predictive accuracy. However, taking a closer look at the single numbers, we can observe that the higher accuracy is not accompanied by a smaller forgetting value. Thus, our results indicate that the most representative samples are beneficial in improving the predictive performance of the learning model and in reducing considerably CF.
The comparison among the different uncertainty strategies suggests that BI outperforms the other metrics in most cases. While the difference among the accuracy values is less pronounced, BI stands out in its effectiveness at mitigating CF. Given the reduced information loss compared to the most popular uncertainty metrics, we think that BI is more sensitive to changes in the predictive space and thus able to detect changes when subtle differences are present. We can notice that MC delivers the smallest forgetting value (F) in most of the cases. However, this is because MC performs poorly in the predictive task (see Figure 3) and considering how forgetting is computed, the resulting value is inevitably smaller. The poor performance of MC compared to other strategies is probably given by the online nature of the setting since, as reported in the original paper, dropout takes longer to converge (Gal & Ghahramani, 2016). Thus, a single-epoch setup may be insufficient for training convergence. All these findings are evident in Figure 3 where we can observe that BI is able to maintain competitive predictive performance on all tasks at the end of the training compared to other strategies.
Figure 3: Average task accuracy at the end of the learning procedure for different strategies and datasets. One can observe how the proposed strategy based on the BI perform consistently across tasks and helps maintaining competitive performance on old tasks and thus reducing catastrophic forgetting.
Figure 4: Class-specific samples stored in the memory according to different sorting strategies: bottom-k selects prototypical examples of the considered class facilitating the recollection of the class characteristics and reducing CF; top-k selects the most uncertain and hard samples making difficult to learn clear discriminative patterns and to reduce CF.